OpenStudy (anonymous):

Which expression is the simplified form of the square root of z^7?

mathslover (mathslover):

First of all , I will introduce you to a property that we are going to use here. \(a^b \times a^c = a^{b+c}\)

mathslover (mathslover):

So we are going to write : \(z^7\) as in the form of \(z^4\) ,\(z^2 \) , \(z\) .

mathslover (mathslover):

Can you tell me how can we write \(z^7\) in a more simplified form.

OpenStudy (anonymous):

im a little confused,

mathslover (mathslover):

Ok! no problem. Can I write : \(z^7 \) as \(z^{1 + 4 + 2}\) ?

OpenStudy (anonymous):

yes

mathslover (mathslover):

So as mentioned above : \(a^{b+c} = a^b \times a^c\) Similarly : \(a^{b+c+d} = a^b \times a^c \times a^d\) Can you now tell me what I can write : \(z^{1+2+4}\) as ? Hint : z = a, 1 = b , 2 = c , 4 = d.

OpenStudy (anonymous):

\[z ^{1}*z ^{2}*z ^{4}\]

OpenStudy (anonymous):

is that right?

mathslover (mathslover):

Yeah !!! Very right. So now we have : \(\sqrt{z^7} = \sqrt{z^1 \times z^2 \times z^4}\) Right?

OpenStudy (anonymous):

right!!

OpenStudy (anonymous):

so since this exponent is odd, do i subtract 1 and leave that one letter inside the radical???

OpenStudy (agent0smith):

\[\Large \sqrt{z^7} = (z^7)^{\frac{ 1 }{ 2}} = \] \[\LARGE (z^1 z^6)^{\frac{ 1 }{ 2 }} = z^{\frac{ 1 }{2 }} z^{\frac{ 6}{ 2 }}\] The last step is just using the fact that \[\Large (ab)^x = a^x b^x\] Now you can simplify. Depending on your class, you may need absolute value signs on the z which will now be outside the square root.

OpenStudy (phi):

can you simplify \( \sqrt{z^6} \) ? because \( \sqrt{z^7} = \sqrt{z^6}\sqrt{z^1} \)

OpenStudy (phi):

or use square root is the same as using the 1/2 exponent.... as agent posted

OpenStudy (anonymous):

so my answer would be \[z ^{3}\sqrt{z}?\]

OpenStudy (phi):

yes. notice that if you simplify \[ z^{\frac{6}{2}} z^{\frac{1}{2}}\] you get \[ z^3 z^{\frac{1}{2}}\] which can be written as \[ z^3 \sqrt{z} \]

OpenStudy (anonymous):

yay i get it!! thank you all for ur help!!

OpenStudy (anonymous):

phi will u give agent0smith a medal (: they deserve one too!